find the midpoint
(-2 7/8, 5/8), (-3 1/4, 1/2)
1
Expert's answer
2020-08-27T17:43:18-0400

How do you find a number halfway between two numbers? You take their average. For instance, what's halfway between "10" and "20"? It's "\frac{10+20}2=15".


You can use the same argument for two points. Since we need to find the midpoint of two 2D points, just take the average of their corresponding x and y coordinates.

Thus, x coordinate of the midpoint

"=\frac{(-2 \frac78)+(-3 \frac14)}2=\frac{(-\frac{23}8)+(-\frac{13}4)}2=-\frac{49}{16}"

y coordinate of the midpoint

"=\frac{\frac58+\frac12}2=\frac{9}{16}"

Thus, the midpoint of the two points is "\left(-\frac{49}{16},\frac9{16}\right)".